3/4(1/2+2/5/1/3)-1/5*3/8
1
Expert's answer
2012-10-15T07:43:36-0400
3/4(1/2+2/5/1/3)-1/5*3/8=3/4(1/2+2/15)-3/40=3/8+6/60-3/40=3/8+1/10-3/40=(15+4-3)/40=16/40=4/10=2/5